# CIVIL: Causal and Intuitive Visual Imitation Learning

### v2

@@ Abstract
Today’s robots attempt to learn new tasks by imitating human examples. These robots watch the human complete the task, and then try to match the actions taken by the human expert. However, this standard approach to visual imitation learning is fundamentally limited: the robot observes what the human does, but not why the human chooses those behaviors. Without understanding the features that factor into the human’s decisions, robot learners often misinterpret the human’s examples (e.g., the robot incorrectly thinks the human picked up a coffee cup because of the color of clutter in the background). In practice, this results in causal confusion, inefficient learning, and robot policies that fail when the environment changes. We therefore propose a shift in perspective: instead of asking human teachers just to show what actions the robot should take, we also enable humans to intuitively indicate why they made those decisions. Under this paradigm human teachers can attach markers to task-relevant objects and use natural language prompts to describe important features. Our proposed algorithm, CIVIL, leverages this augmented demonstration data to filter the robot’s visual observations and extract a feature representation that causally informs human actions. CIVIL then applies these causal features to train a transformer-based policy that — when tested on the robot — is able to emulate human behaviors without being confused by visual distractors or irrelevant items. Our simulations and real-world experiments demonstrate that robots trained with CIVIL learn both what actions to take and why to take those actions, resulting in better performance than state-of-the-art baselines. From the human’s perspective, our user study reveals that this new training paradigm actually reduces the total time required for the robot to learn the task, and also improves the robot’s performance in previously unseen scenarios. See videos at our project website: https://civil2025.github.io
I Introduction
@@
Imitation learning enables robots to learn new tasks by emulating the actions of a human expert. Consider a human teaching a robot arm to serve coffee (as shown in Figure 1). The human guides the robot through different states of the task, including picking a cup off the kitchen counter and placing it under the coffee machine. To learn the task, the robot observes the scene with an onboard camera and records the actions demonstrated by the human teacher. But these demonstrations only tell the robot what it should do, leaving the robot to figure out why it should perform these actions (i.e., what aspects of the environment factored into the human’s decisions).
@@
Understanding the reasoning behind the human’s actions is critical for adapting to new situations. For example, humans know that the coffee cup’s position affects how it should be grasped; if the cup moves, humans will change their actions to reach its new position. However, it is difficult for robots to infer this underlying goal solely from the demonstrated actions because their visual observations often contain excess information — along with the cup, the robot’s camera also sees other utensils and appliances on the kitchen counter. These irrelevant details can create causal confusion when they are correlated with the human’s actions [1]. For instance, if the cup is always next to a bowl during the demonstrations, the robot may not understand the human’s motive; should it reach for the cup or go to a position beside the bowl?
@@
Figure 1: Human teaching a robot arm to prepare a cup of coffee. The robot must learn to grasp the cup and place it under a coffee machine based on visual observations. Within traditional approaches the human demonstrates what actions to take, and the robot learns to emulate these demonstrated actions. However, this approach is inefficient because the robot is not taught why the human chooses a specific behavior. Without causal information the robot can misinterpret the human: for instance, if a bowl is always placed to the left of the cup during the demonstrations, the robot might learn to go beside the bowl instead of go to the cup. We hypothesize that robots can learn more efficiently when the human communicates the features behind their decisions (i.e., why they are choosing the actions they demonstrate). CIVIL shifts imitation learning towards holistic demonstrations with physical markers and natural language instructions.
@@
Existing research has focused on enabling robots to resolve this confusion on their own by making assumptions about task-relevant information. Current imitation learning methods try to extract the relevant details from the robot’s observations by augmenting the data with random transformations [2, 3], identifying known objects in the scene [4], or using vision-language models pretrained on large datasets [5, 6]. While these approaches help robots adapt their actions to expected variations of the task, they require a significant amount of data to truly uncover the human’s reasoning. For example, when we experimentally applied these baselines to the task in Figure 1, we found that the robot may incorrectly learn to focus on the bowl (instead of the coffee cup) because of spurious correlations in the training data. This leads to robots that cannot make coffee when the bowl is removed.
@@
To address this fundamental limitation we here re-frame the process of learning from human demonstrations. Rather than expecting robots to infer the correct causality based solely on human actions, we now extend imitation learning so that human teachers can intuitively reveal what actions to take and why to take those actions. Our hypothesis is:
@@
Robots can learn tasks more effectively when the human provides a smaller number of demonstrations while communicating the key features behind their actions.
@@
We apply this hypothesis to create interfaces that humans can leverage to convey causality during their demonstrations. Specifically, we use a combination of physical markers and language instructions to give context to human demonstrations. Human teachers place markers in the environment to highlight relevant objects, positions, and interactions that inform their actions (i.e., the human in Figure 1 might mark the coffee cup and coffee machine). Similarly, the human can provide natural language utterances to explain what they are doing or what they are focusing on during their demonstration (i.e., “pick up the cup”). The robot learner collects the demonstrated state and actions — as in traditional approaches — along with the new marker positions and language prompts.
@@
These augmented demonstrations provide the robot with a more holistic understanding of the task and supplement its learning in two ways. First, the robot leverages the marker and language cues to filter its extraneous observations and extract a low-dimensional feature representation that encodes human reasoning. Second, the robot learns a policy that maps these causal features to the demonstrated actions while remaining robust to unintended correlations and irrelevant visual data. We refer to our resulting algorithm as:
CIVIL: Causal and Intuitive Visual Imitation Learning
Using CIVIL, humans can teach robots the true intentions (why) behind their actions (what). Our results reveal that users perceive this immersive teaching protocol to be more intuitive and natural (i.e., how humans would teach other humans). We also emphasize that gathering the additional marker and language data does not increase the overall teaching burden: instead, we find that users require fewer demonstrations and less total time to train the robot, and the resulting robot policy is more robust to new scenarios.
This work is a step towards robots that are able to correctly understand and perform tasks based on a few human demonstrations. Overall, we make the following contributions:
@@
Comparing to State-of-the-Art Alternatives. We evaluate robots that act on the human-supervised features of CIVIL against multiple state-of-the-art baselines that let robots derive causality through self-supervision [2], object detection [4], and pre-trained vision-language models [7]. Our experiments include simulations in CALVIN [8], a benchmark for learning manipulation tasks, as well as real-world experiments with a Franka Emika Panda robot arm. Robots trained on CIVIL were more successful in performing the tasks than the baselines, especially when tested on unseen task instances.
@@ II Related Work
II-A Visual Imitation Learning
@@
When imitating humans, the robot learns a policy that maps its observations to the actions demonstrated by a human expert. We expect robots to learn this expert policy from a few demonstrations and then transfer it to other, potentially unseen variations of the task [9, 10]. But when the observations are high-dimensional and contain extraneous information, it can be difficult for robots to infer which parts of these observations actually affect the task performance [1]. For example, the robot in Figure 1 may not know which objects to focus on when making coffee on a cluttered kitchen counter.
@@
To resolve this confusion, robots can encode their observations into a low-dimensional feature representation that only retains essential information — such as the position of the cup — and ignores irrelevant details like lighting changes and background objects [11]. Existing approaches let robots derive these features on their own by making assumptions about the extraneous aspects [2, 12, 13, 5], or simply focusing on the known objects in the scene [14, 4, 15].
@@
For instance, the robot can generate alternative views of the images taken from its camera by applying transformations like color distortions and random cropping [16, 17, 18], and then train an encoder to map these transformed views into the same features as the original, unmodified images. This helps the robot learn a feature representation that is invariant to noisy transformations. Alternatively, the robot can use existing vision models to detect known objects in its view and train its policy on features derived from the segmented images of these objects [4]. This approach encourages the robot to disregard any background details like the appearance of the kitchen counter or the lighting of the room.
@@
While these unsupervised approaches make the robot robust to distractors like lighting and background, they rely on the robot to implicitly infer the relevant features (e.g., the cup’s position) from human demonstrations. This slows the learning process — humans need to provide demonstrations in diverse scenarios to facilitate causal inference [19] — and can also be counterproductive when the assumed variations deviate from the human’s reasoning [17]. For example, if a user wants the robot to interact with objects of a specific color or focus on some background cues, training with images that vary in color or exclude the background can further confuse the robot.
@@
Hence, in this work we enable humans to explicitly convey their underlying representations to the robot. We anticipate that communicating the reasoning behind human actions will mitigate causal confusion and accelerate learning. Accordingly, we next discuss prior works that have explored how humans can intuitively reveal their intentions to robot learners.
II-B Learning Human Representations
@@
Robots can learn more efficiently and generalize better to unseen scenarios when their representations are aligned with human reasoning [20]. To achieve this alignment, humans need to share further insights into their decision-making while demonstrating the task. Earlier works have proposed obtaining representations by asking humans to select the task-relevant factors from a pre-defined list [21, 22, 23], label the features for examples in the training data [24, 25], and provide demonstrations that trace the gradient of a relevant feature [26]. However, these approaches are either cognitively demanding because users find it difficult to quantify feature values [27], or physically taxing due to the need for additional feature-specific demonstrations. To feasibly obtain this information in practice, it is important to leverage natural and intuitive communication channels that can be seamlessly integrated into the robot’s training process [28]. Therefore, recent work has focused on pairing demonstrations with natural language prompts [5, 29, 30, 31, 32, 33, 34, 6], and introducing intuitive sensors and interfaces to collect additional human inputs [35, 36, 37, 38, 39, 40, 41, 42, 43].
@@
Humans can organically explain their actions using natural language. For example, when demonstrating how to make coffee, users may say “pick up the cup” and then “place it under the coffee machine.” Prior works have shown that robots can utilize these prompts to improve their representations in multiple ways. Many previous approaches encode language descriptions into feature vectors and pair them with visual features to provide more context for the robot’s policy [29, 30, 31, 32, 33, 34]. Some works use language to supervise how features are extracted from robot observations by using contrastive learning, as in CLIP [7], or by conditioning their visual encoder [3]. Lastly, instead of learning the features from scratch, we can take pretrained vision-language models and fine-tune them on demonstrations of the task [5, 6]. In our work, instead of using language to contextualize the robot’s features or policy, we leverage language to filter the robot’s observations — highlighting relevant objects in the scene and removing irrelevant details that can confuse the learner.
@@
Although humans can explain parts of their thinking using natural language, not every aspect of a task can be easily put into words, e.g., subconscious visual cues or complex motion constraints. Such details can be communicated more intuitively through specialized instruments. For instance, humans can cheaply convey rich motion information using hand-held grasping tools [35, 36], optical trackers [37], and wearable tactile gloves [38]. Humans can also utilize augmented reality interfaces to specify keyframes and motion constraints [39, 40]. Alternatively, the robot can track human gaze and focus on the same regions of its observations as the expert user [41, 42]. Most relevant to our approach are prior works that either use Bluetooth sensors to locate relevant objects in the environment [44], or introduce an interface for humans to mark these objects on images of the scene [45].
@@ III Problem Statement
Features. The robot’s observations are high-dimensional and contain both relevant information for learning the task and extraneous details that should be ignored. For instance, along with the cup that we want the robot to grasp, it could also see a bowl and other kitchen appliances on the counter. The robot does not know which parts of these observations are relevant a priori. We represent the task-relevant information as a compact feature vector $\phi^{*}\in\mathbb{R}^{d}$, where the feature dimension $d$ is less than the dimensionality $n$ of the observations. In our example, $\phi^{*}$ contains the cup’s position and orientation but excludes the bowl and other irrelevant items on the kitchen counter.
@@
In the above $\theta^{*}$ are the true parameters of the policy that the human wants to teach the robot. Intuitively, the policy parameters $\theta^{*}$ dictate what actions the human will take and the features $\phi^{*}$ determine why the human chooses that action for a given robot state and observation.
@@
III-A Using Low-Dimensional Features to Accelerate Learning

### v3

@@ Abstract
Today’s robots attempt to learn new tasks by imitating human examples. These robots watch the human complete the task, and then try to match the actions taken by the human expert. However, this standard approach to visual imitation learning is fundamentally limited: the robot observes what the human does, but not why the human chooses those behaviors. Without understanding which features of the system or environment factor into the human’s decisions, robot learners often misinterpret the human’s examples (e.g., the robot incorrectly thinks the human picked up a coffee cup because of the color of clutter in the background). In practice, this results in causal confusion, inefficient learning, and robot policies that fail when the environment changes. We therefore propose a shift in perspective: instead of asking human teachers just to show what actions the robot should take, we also enable humans to intuitively indicate why they made those decisions (i.e., what features are critical for the desired task). Under our paradigm human teachers attach markers to task-relevant objects and use natural language prompts to describe their state representation. Our proposed algorithm, CIVIL, leverages this augmented demonstration data to filter the robot’s visual observations and extract a feature representation that aligns with the human teacher. CIVIL then applies these causal features to train a transformer-based policy that — when tested on the robot — is able to emulate human behaviors without being confused by visual distractors or irrelevant items. Our simulations and real-world experiments demonstrate that robots trained with CIVIL learn both what actions to take and why to take those actions, resulting in better performance than state-of-the-art baselines. From the human’s perspective, our user study reveals that this new training paradigm actually reduces the total time required for the robot to learn the task, and also improves the robot’s performance in previously unseen scenarios. See videos at our project website: https://civil2025.github.io
journal: Autonomous Robots
1 Introduction
@@
Imitation learning enables robots to learn new tasks by emulating the actions of a human expert. Consider a human teaching their robot arm to serve coffee (as shown in Figure 1). The human guides the robot through different stages of the task, including picking a cup off the kitchen counter and placing it under the coffee machine. To learn this task, the robot observes the scene with an onboard camera and records the actions demonstrated by the human teacher. But these visual demonstrations only show the robot what it should do, leaving the robot to figure out why it should perform these actions (i.e., what aspects of the system and environment states factored into the human’s decisions).
@@
Understanding the state representation behind the human’s actions is critical for adapting to new situations. For example, humans know that the coffee cup’s position affects how it should be grasped; if the cup moves, humans will change their actions to match its new configuration. However, it is difficult for robots to infer this underlying feature solely from the demonstrated actions because their visual observations often contain excess information — along with the cup, the robot’s camera also sees other utensils and appliances on the kitchen counter. These irrelevant details can create causal confusion when they are correlated with the human’s actions de2019causal . For instance, if the cup is always next to a bowl during the demonstrations, the robot may not understand the human’s motive; should it reach for the cup or go to a position beside the bowl? Put another way — what features of the observed state are connected to the task, and which features are irrelevant clutter or spurious correlations?
@@
Figure 1: Human teaching a robot arm to prepare a cup of coffee. The robot must learn to grasp the cup and place it under a coffee machine based on visual observations. Within traditional approaches the human demonstrates what actions to take, and the robot learns to emulate these demonstrated actions. However, this approach is inefficient because the robot is not taught why the human chooses a specific behavior (i.e., what features of the environment factored into the human’s decisions). Without this causal information that links features to actions the robot can misinterpret the human: for instance, if a bowl is always placed to the left of the cup during the demonstrations, the robot might learn to go beside the bowl instead of go to the cup. We hypothesize that robots can learn more efficient and robust control policies when the human teacher communicates the features behind their decisions (i.e., why they are choosing the actions they demonstrate). CIVIL shifts imitation learning towards holistic demonstrations with physical markers and natural language instructions.
@@
Existing research has focused on enabling robots to resolve this confusion on their own by making assumptions about task-relevant information. Current imitation learning methods try to extract the relevant details from the robot’s observations by augmenting the data with random transformations pari2021surprising ; karamcheti2023language , identifying known objects in the scene zhu2023viola , or using vision-language models pretrained on large datasets sharma2023lossless ; yang2024robot . While these approaches help robots adapt their actions to expected variations of the task, they require a significant amount of data to truly uncover the human’s reasoning. For example, when we experimentally applied these baselines to the task in Figure 1, we found that the robot may incorrectly learn to focus on the bowl (instead of the coffee cup) because of misleading correlations in the training data. This leads to robots that cannot make coffee when the bowl is removed.
@@
To address this fundamental limitation we here re-frame the process of learning from human demonstrations. Rather than expecting robots to infer the correct causality based solely on human actions, we now extend imitation learning so that human teachers can intuitively reveal what actions to take and why to take those actions (i.e., what environment features guided the human’s demonstrations). Our hypothesis is:
@@
Robots can learn more effectively when the human provides a smaller number of demonstrations while communicating the key features behind their actions.
@@
We apply this hypothesis to create interfaces that humans can leverage to convey state representations during their demonstrations. Specifically, we use a combination of physical markers and language instructions to give context to human demonstrations. Human teachers place markers in the environment to highlight relevant objects, positions, and interactions that inform their actions (i.e., the human in Figure 1 might mark the coffee cup and coffee machine). Similarly, the human can provide natural language utterances to explain what they are doing or what they are focusing on during their demonstration (i.e., “pick up the cup”). The robot learner collects the demonstrated state and actions — as in traditional approaches — along with the new marker positions and language prompts.
@@
These augmented demonstrations provide the robot with a more holistic understanding of the task and supplement its learning in two ways. First, the robot leverages the marker and language cues to filter its extraneous observations and extract a low-dimensional feature representation that encodes human reasoning. Second, the robot learns a policy that maps these causal features to the demonstrated actions while remaining robust to unintended correlations and irrelevant visual data. We refer to our resulting algorithm as CIVIL: Causal and Intuitive Visual Imitation Learning. Using CIVIL, humans can provide the robot with labeled data about the system and environment features (why) that guided their demonstrationed actions (what). Our results reveal that users perceive this immersive teaching protocol to be more intuitive and natural (i.e., how humans would teach other humans). We also emphasize that gathering the additional marker and language data does not increase the overall teaching burden: instead, we find that users require fewer demonstrations and less total time to train the robot, and the resulting robot policy is more robust to new scenarios.
This work is a step towards robots that are able to correctly understand and perform tasks based on a few human demonstrations. Overall, we make the following contributions [^1] A preliminary version of this work was published at the IEEE International Conference on Intelligent Robots and Systems sanchez2024recon . As compared to sanchez2024recon , this manuscript i) provides theoretical analysis justifying our proposed demonstration paradigm, ii) develops an end-to-end vision and language network that extracts causal features from human guidance, and iii) compares CIVIL to state-of-the-art alternatives while evaluating how humans interact with CIVIL.:
@@
Comparing to State-of-the-Art Alternatives. We compare robots that act on the human-supervised features of CIVIL against multiple state-of-the-art baselines that let robots derive causality through self-supervision pari2021surprising , object detection zhu2023viola , and pre-trained vision-language models radford2021learning . Our experiments include simulations in CALVIN mees2022calvin , a benchmark for learning manipulation tasks, as well as real-world experiments with Franka robot arms. Robots trained on CIVIL are more successful in performing the tasks than the baselines, especially when tested on unseen task instances.
@@ 2 Related Work
2.1 Visual Imitation Learning
@@
When imitating humans, the robot learns a policy that maps its observations to the actions demonstrated by a human expert. We expect robots to learn this expert policy from a few demonstrations and then transfer it to other, potentially unseen variations of the task osa2018algorithmic ; mandi2022towards . But when the observations are high-dimensional and contain extraneous information, it can be difficult for robots to infer which parts of these observations actually affect the task performance de2019causal . For example, the robot in Figure 1 may not know which objects to focus on when making coffee on a cluttered kitchen counter.
@@
To resolve this confusion, robots can encode their observations into a low-dimensional feature representation that only retains essential information — such as the position of the cup — and ignores irrelevant details like lighting changes and background objects xie2024decomposing . Existing approaches let robots derive these features on their own by making assumptions about the extraneous aspects pari2021surprising ; radosavovic2023real ; pfrommer2023initial ; sharma2023lossless , or simply focusing on the known objects in the scene qin2022dexmv ; zhu2023viola ; jonnavittula2025view .
@@
For instance, the robot can generate alternative views of the images taken from its camera by applying transformations like color distortions and random cropping chen2020simple ; grill2020bootstrap ; he2022masked , and then train an encoder to map these transformed views into the same features as the original, unmodified images. This helps the robot learn a feature representation that is invariant to noisy transformations. Alternatively, the robot can use existing vision models to detect known objects in its view and train its policy on features derived from the segmented images of these objects zhu2023viola . This approach encourages the robot to disregard background details like the appearance of the kitchen or the lighting of the room.
@@
While these unsupervised approaches make the robot robust to distractors like lighting and background, they rely on the robot to implicitly infer the relevant features (e.g., the cup’s position) from human demonstrations. This slows the learning process — humans need to provide demonstrations in diverse scenarios to facilitate causal inference bica2021invariant — and can also be counterproductive when the assumed variations deviate from the human’s reasoning grill2020bootstrap . For example, if a user wants the robot to interact with objects of a specific color or focus on some background cues, training with images that vary in color or exclude the background can further confuse the robot.
@@
Hence, in this work we enable humans to explicitly convey their underlying feature representations to the robot. We anticipate that communicating the reasoning behind human actions will mitigate causal confusion and accelerate learning. Accordingly, we next discuss prior works that have explored how humans can intuitively reveal their intentions to robot learners.
2.2 Learning Human Representations
@@
Robots can learn more efficiently and generalize better to unseen scenarios when their representations are aligned with human reasoning bobu2024aligning . To achieve this alignment, humans need to share further insights into their decision-making while demonstrating the task. Earlier works have proposed obtaining representations by asking humans to select the task-relevant factors from a pre-defined list cakmak2012designing ; basu2018learning ; nemlekar2023transfer , label the features for examples in the training data sripathy2022teaching ; schrum2024maveric , and provide demonstrations that trace the gradient of a relevant feature bobu2022inducing . However, these approaches are either cognitively demanding because users find it difficult to quantify feature values koppol2021interaction , or physically taxing due to the need for additional feature-specific demonstrations. To feasibly obtain this information in practice, it is important to leverage natural and intuitive communication channels that can be seamlessly integrated into the robot’s training process habibian2024survey . Therefore, recent work has focused on pairing demonstrations with natural language prompts sharma2023lossless ; mees2022matters ; jang2022bc ; yu2023using ; ma2024actra ; brohan2023rt ; team2024octo ; yang2024robot , and introducing intuitive sensors and interfaces to collect additional human inputs song2020grasping ; young2021visual ; song2023data ; wei2024wearable ; quintero2018robot ; luebbers2021arc ; saran2021efficiently ; biswas2024gaze ; nemlekar2024pecan .
@@
Humans can organically explain their actions using natural language. For example, when demonstrating how to make coffee, users may say “pick up the cup” and then “place it under the coffee machine.” Prior works have shown that robots can utilize these prompts to improve their representations in multiple ways. Many previous approaches encode language descriptions into feature vectors and pair them with visual features to provide more context for the robot’s policy mees2022matters ; jang2022bc ; yu2023using ; ma2024actra ; brohan2023rt ; team2024octo . Some works use language to supervise how features are extracted from robot observations by using contrastive learning, as in CLIP radford2021learning , or by conditioning their visual encoder karamcheti2023language . Lastly, instead of learning the features from scratch, we can take pretrained vision-language models and fine-tune them on demonstrations of the task sharma2023lossless ; yang2024robot . In our work, instead of using language to contextualize the robot’s features or policy, we leverage language to filter the robot’s observations — highlighting relevant objects in the scene and removing irrelevant details that can confuse the learner.
@@
Although humans can explain parts of their thinking using natural language, not every aspect of a task can be easily put into words, e.g., subconscious visual cues or complex motion constraints. Such details can be communicated more intuitively through specialized instruments. For instance, humans can cheaply convey rich motion information using hand-held grasping tools song2020grasping ; young2021visual , optical trackers song2023data , and wearable tactile gloves wei2024wearable . Humans can also utilize augmented reality interfaces to specify keyframes and motion constraints quintero2018robot ; luebbers2021arc . Alternatively, the robot can track human gaze and focus on the same regions of its observations as the expert user saran2021efficiently ; biswas2024gaze . We explored this option in our preliminary work sanchez2024recon , where humans used Bluetooth sensors to locate relevant objects in the environment; similarly, zhu2023learning introduced an interface for humans to mark these objects on images of the scene zhu2023learning . However, we find that physical markers alone are not sufficient to capture critical features. These markers might indicate where the robot should focus (e.g., “look at the light bulb”), but not what aspects to focus on (e.g., “check if the light is on or off”).
@@ 3 Problem Statement
Features. The robot’s observations are high-dimensional and contain both relevant information for learning the task and extraneous details that should be ignored. For instance, along with the cup that we want the robot to grasp, it could also see a bowl and other kitchen appliances on the counter. The robot does not know which parts of these observations are relevant a priori. We represent the task-relevant information as a compact feature vector $\phi^{*}\in\mathbb{R}^{d}$, where the feature dimension $d$ is less than the dimensionality $n$ of the observations. In our example, $\phi^{*}$ contains the cup’s position and orientation but excludes the bowl and other irrelevant items on the kitchen counter. Our work focuses on extracting the relevant features from the human’s demonstrations so that we can map high-dimensional states into compact feature vectors.
@@
In the above $\theta^{*}$ are the true parameters of the policy that the human wants to teach the robot. Intuitively, the policy parameters $\theta^{*}$ dictate what actions the human will take and the features $\phi^{*}$ determine why the human chooses that action for a given robot state and observation (i.e., $\phi^{*}$ provides a feature state representation for the desired task).
@@
3.1 Using Low-Dimensional Features to Accelerate Learning